(a) If the molar solubility of Mg3(PO4)2 at 25 oC is 6.26e-06 mol/L, what is the Ksp at this temperature?

Ksp =




(b) It is found that 6.35e-07 g of Cu3(PO4)2 dissolves per 100 mL of aqueous solution at 25 oC. Calculate the solubility-product constant for Cu3(PO4)2.

Ksp =




(c) The Ksp of PbBr2 at 25 oC is 6.60e-06. What is the molar solubility of PbBr2?

solubility =  mol/L
Answer #1

a.

the molar solubility of Mg3(PO4)2 at 25 oC is 6.26*10^-6 mol/L

Mg3PO4)2(s) ---------> 3Mg^2+ (aq) + 2PO4^3- (aq)

                                     3s                      2s

Ksp     = [Mg^2+]^3[PO4^3-]^2

          = (3s)^3(2s)^2

         = 108s^5

          = 108*(6.26*10^-6)^5

          = 1.04*10^-24

b.

molarity   = W*1000/G.M.Wt* volume in ml

              = 6.35*10^-7*1000/(380.58*100)

               = 1.7*10^-8M

Cu3(PO4)2(s) ------------> 3Cu^2+ (aq) + 2PO4^3-(aq)

                                            3s                    2s

Ksp   = [Cu^2+]^3[PO4^3-]^2

        = (3s)^3(2s)^2

       = 108s^5

       = 108(1.7*10^-8)^5

         = 1.135*10^-38

c. PbBr2(s) ------------. Pb^2+ (aq) + 2Br^- (aq)

                                   s                      2s

Ksp    = [Pb^2+][Br^-]^2

6.6*10^-6   = s(2s)^2

6.6*10^-6   = 4s^3

s   = 0.0118

solubility of PbBr2 = 1.18*10^2mole/L